There are 50 apple trees in an orchard. Each tree produces 800 apples. For each additional tree planted in the orchard, the
output per tree drops by 10 apples. How many trees should be added to the existing orchard in order to maximize the total
output of trees ?

Grade:12

5 Answers

bharat bajaj IIT Delhi
10 years ago
The output for x no. of trees is then : (800 - 10n)(50+n)
Hence, O = 300n - 10n^2 + 40000
Differentiate it :
300 - 20n = 0
hence, n = 15
Hence, 15 trees should be added.
Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi
sidduchinnu
10 years ago
x no. of trees is then : (800 - 10n)(50+n) = 300n - 10n^2 + 4000 differentiate it : 300 - 20n = 0 n = 15 there fore, 15 trees should be added.
Sher Mohammad IIT Delhi
10 years ago
let number of additional tree is n , now total trees =50+n
apple production per tree =800-10n
total production =(50+n)(800-10n)
= -(10n^2-300n-40000)
f'(n)= 20n-300=0
n=15
sher mohammad
b.tech, iit delhi

Sher Mohammad IIT Delhi
10 years ago
let number of additional tree is n , now total trees =50+n
apple production per tree =800-10n
total production =(50+n)(800-10n)
= -(10n^2-300n-40000)
f'(n)= 20n-300=0
n=15
sher mohammad
b.tech, iit delhi
Arun Kumar IIT Delhi
9 years ago


Arun Kumar
IIT Delhi
Askiitians Faculty